# VietLyrics: A Large-Scale Dataset and Models for Vietnamese Automatic Lyrics Transcription

### v1

@@ Abstract
This is the author’s accepted version of the paper: Nguyen Quoc Anh, Bernard Cheng, and Kelvin Soh, “VietLyrics: A Large-Scale Dataset and Models for Vietnamese Automatic Lyrics Transcription,” 2025 Asia Pacific Signal and Information Processing Association Annual Summit and Conference (APSIPA ASC), 2025. The final version is available on IEEE Xplore via DOI: https://doi.org/10.xxxx/APSIPAASC.2025.xxxxx.
@@ 1 Introduction
  • The Whisper [1] architecture was further fine-tuned using the VietLyrics dataset, significantly improving performance over existing multilingual ALT/ASR systems. Three model variants of different sizes are released to support further research and development in Vietnamese lyrics transcription.
@@ 2 Related Works
DALI-full [2] 30 (mostly EN) 5,358
DALI-train [2] 1 (EN) 3,913
MulJam [3] 6 (EN, FR, DE, ES, IT, RU) 6,031 (381.9h)
JamendoLyrics [4] 4 (EN, ES, DE, FR) 80
DSing30 [5] 1 (EN) 4,324 (149.1h)
MUSDB18 [6] 1 (EN) 150 (10h)
@@ 2.1 Dataset Scarcity
Table 1 presents a non-exhaustive list of open-source song-lyric datasets. There is a lack of large-scale non-English song datasets in general, due to copyright law. These datasets circumvent this issue by collecting music exclusively under Creative Commons licenses or in the Public Domain [2, 4, 6], leveraging AI-generated lyrics [3], and incorporating karaoke vocals voluntarily shared by app users [5]. None of these include Vietnamese, based on our findings, which leads to the lack of Vietnamese ALT models.
@@ 2.2 Limitations in Existing Solutions
Among existing ALT frameworks, LyricWhiz [3] is the current SOTA for multilingual ALT, leveraging both Whisper for transcription and GPT-4 for refining outputs. Whisper serves as the “ear” by converting audio into text, while GPT-4 functions as the “brain”, selecting the most accurate lyrics. While LyricWhiz has demonstrated strong performance across various prominent languages, its effectiveness on Vietnamese lyrics is limited. The replicated baseline model struggles with Vietnamese audio, as shown in the qualitative results in Table 2. Note that other solutions also suffer from similar errors. Most errors stem from the inherent complexity of Vietnamese phonetics and the fact that Whisper, being primarily an ASR model, is not optimized for ALT tasks.
@@ 3.1.1 Data Collection & Processing
Under Article 25, Clause 1(a) of the Vietnamese Intellectual Property Law (amended June 16, 2022), copyrighted musical works may be reproduced for academic and research purposes without permission or fees, provided the original creators are credited [7]. To respect copyright, we also scraped 2 other lyrics websites [8, 9] for songwriters and artists for proper attributions, to the best of our abilities. We publicly release all metadata (including song names, links, and author attributions) and scraping code for academic research, allowing future studies to recreate the dataset, train, and benchmark for automatic transcription or other tasks.
@@ 3.1.2 Data Analysis
The dataset includes songs from over 4,000 unique artists, lasting on average 4.6 minutes/song. The average singing speed is found to be 90.1 Words Per Minute (WPM), less than half the Vietnamese spoken rate of 190 WPM [10]. The dataset includes various genres, as shown in Fig. 1. However, metadata extraction from zingmp3 alone resulted in many songs lacking genre information. Since Vietnamese music research lacks standardized genre classification tools or an agreed-upon genre ontology, Contrastive Learning of Musical Representations (CLMR) [11] was used to infer the top 50 MagnaTagATune tags [12]. The result was of weak alignment, with the 95th percentile confidence score reaching only 0.22. This suggests that this collection of Vietnamese music does not fit well within existing Western music’s tag ontology, highlighting the need for further research in Vietnamese music genre classification.
@@
Vietnamese dialects (which include both accents and variations in word choices, idioms, sentence structure, etc.) differ significantly across regions, a difference even more pronounced in singing, making accent classification valuable for music analysis. A pre-trained Wav2Vec2-base model [13, 14] fine-tuned on Vietnamese Speech corpora [15, 16] was used to predict gender [17] and dialect [18].
We extracted and ran inference on three 10-second vocal-only segments (isolated using Demucs v3 [19]) and determined the final classification through majority voting. For gender classification specifically, a “both” category was introduced for cases where at least one segment had a confidence score between 0.4 and 0.6, indicating ambiguity due to a unisex voice, absence of vocals, or a mix of male and female vocals. As shown in Fig. 2, the results indicate a distribution skewed towards Northern dialects, while Central dialects are underrepresented. Male-sung songs were more prevalent than those classified as “female” or “both”. This data imbalance is hard to address, as gender and accent metadata were unavailable during data collection. (Note that this observed skew can also stem from Wav2Vec2 model bias.)
@@ 3.2 Modeling
Since the current SOTA LyricWhiz model does not effectively handle Vietnamese songs, our approach is to fine-tune an ASR model end-to-end to serve as an ALT system. The Whisper model [1] was chosen as the starting point, as it is renowned for its SOTA performance in multilingual tasks and exceptional fine-tuning capabilities.
Three Whisper model variants of different sizes (small, medium and large-v2) were fine-tuned and made available as checkpoints on Hugging Face [20].
@@ 3.2.1 Data Pre-processing
Audio segmentation of songs is performed using the provided ground truth timestamps from the song lyrics to chunk both the audio and lyrics into 30-second segments. This is required as Whisper processes audio in 30-second chunks [1].
@@ 4.1 Finetuned Models
LyricWhiz was evaluated as a baseline for lyrics transcription, given its state-of-the-art performance in prior studies [3], though it had not been tested on Vietnamese songs. PhoWhisper [21], an ASR model fine-tuned on Vietnamese audio, was also assessed with the expectation of better performance on Vietnamese datasets. On VietLyrics test set, LyricWhiz produced a Word Error Rate (WER) of 49.22%, while PhoWhisper-large achieved 38.3%, shown in Table 3. However, both models fell short of the accuracy needed for effective Automatic Lyrics Transcription.
@@
We also experimented with training and evaluating only on vocals, separated using Demucs v3 [19]. However, the improvement observed after source separation was minimal for both models, with 1.3% WER improvement for whisper-medium and 0.5% WER improvement for whisper-large-v2. Given the negligible benefits and the added complexity of these extra pre-processing steps, the decision was made to exclude these steps from the final approach.
Lower-case evaluation was also found to produce better performance metrics compared to case-sensitive evaluation, with WER decreasing by up to 5% on large, as shown in Table 4. This intuitive result suggests that relying on the LLM-based transcriber for capitalization may not be the optimal approach. Instead, more effective methods for inferring line breaks could involve post-processing techniques, such as detecting silence between singing segments or leveraging the approaches introduced in Cífka et al. [22].
@@
Model
WER
(Case-sensitive)
WER
@@
Since hallucination from non-vocal audio is a major challenge in ASR models, experiments were conducted to suppress <nospeech> tokens. Following the original Whisper paper [1], tokens were filtered when the <nospeech> probability exceeded 0.6 and the log confidence probability was below -1. These experiments were evaluated on 100 randomly sampled songs from the 1,000-song validation set. PhoWhisper was tested under various configurations, with vocals-only tracks extracted using Demucs. As shown in Table 5, for PhoWhisper-small, combining source-separated audio with <nospeech> suppression provided the best performance. However, for Whisper large, testing on raw audio without suppression performed best, with our fine-tuned Whisper significantly outperforming others.
This outcome is consistent with the findings in Cífka et al. [22], where the transcriptions from only vocals are less accurate than those from raw data for large models. This is preferable, as it enables deploying an end-to-end model without dependence on source separation performance. Furthermore, qualitative reviews of our finetuned Whisper’s inferences found no instances of long hallucinations, resulting in the discontinuation of further silence suppression experiments.
@@ 5 Conclusion
We collected VietLyrics, a large-scale diverse dataset of Vietnamese music with lyrics, and fine-tuned the Whisper models [1] to develop an Automatic Lyrics Transcription system. Our best performing model, Whisper-large-v2, achieved a WER of 24.61% for case-sensitive lyrics and 20.52% for lowercase lyrics, demonstrating its effectiveness in transcribing Vietnamese songs of various dialects, genres and genders. We publicly release the VietLyrics dataset [23] and models [24, 25, 26], hoping that this work inspires further research and improvements in Automatic Lyrics Transcription and music computing in general, for the Vietnamese language and music.

### v2

@@ Abstract
This is the author’s accepted version of the paper: Nguyen Quoc Anh, Bernard Cheng, and Kelvin Soh, “VietLyrics: A Large-Scale Dataset and Models for Vietnamese Automatic Lyrics Transcription,” 2025 Asia Pacific Signal and Information Processing Association Annual Summit and Conference (APSIPA ASC), 2025. The final version is available on IEEE Xplore via DOI: https://doi.org/10.1109/APSIPAASC65261.2025.11249339
@@ 1 Introduction
  • The Whisper [undef] architecture was further fine-tuned using the VietLyrics dataset, significantly improving performance over existing multilingual ALT/ASR systems. Three model variants of different sizes are released to support further research and development in Vietnamese lyrics transcription.
@@ 2 Related Works
DALI-full [undefa] 30 (mostly EN) 5,358
DALI-train [undefa] 1 (EN) 3,913
MulJam [undefb] 6 (EN, FR, DE, ES, IT, RU) 6,031 (381.9h)
JamendoLyrics [undefc] 4 (EN, ES, DE, FR) 80
DSing30 [undefd] 1 (EN) 4,324 (149.1h)
MUSDB18 [undefe] 1 (EN) 150 (10h)
@@ 2.1 Dataset Scarcity
Table 1 presents a non-exhaustive list of open-source song-lyric datasets. There is a lack of large-scale non-English song datasets in general, due to copyright law. These datasets circumvent this issue by collecting music exclusively under Creative Commons licenses or in the Public Domain [undefa, undefc, undefe], leveraging AI-generated lyrics [undefb], and incorporating karaoke vocals voluntarily shared by app users [undefd]. None of these include Vietnamese, based on our findings, which leads to the lack of Vietnamese ALT models.
@@ 2.2 Limitations in Existing Solutions
Among existing ALT frameworks, LyricWhiz [undefb] is the current SOTA for multilingual ALT, leveraging both Whisper for transcription and GPT-4 for refining outputs. Whisper serves as the “ear” by converting audio into text, while GPT-4 functions as the “brain”, selecting the most accurate lyrics. While LyricWhiz has demonstrated strong performance across various prominent languages, its effectiveness on Vietnamese lyrics is limited. The replicated baseline model struggles with Vietnamese audio, as shown in the qualitative results in Table 2. Note that other solutions also suffer from similar errors. Most errors stem from the inherent complexity of Vietnamese phonetics and the fact that Whisper, being primarily an ASR model, is not optimized for ALT tasks.
@@ 3.1.1 Data Collection & Processing
Under Article 25, Clause 1(a) of the Vietnamese Intellectual Property Law (amended June 16, 2022), copyrighted musical works may be reproduced for academic and research purposes without permission or fees, provided the original creators are credited [undeff]. To respect copyright, we also scraped 2 other lyrics websites [undefg, undefh] for songwriters and artists for proper attributions, to the best of our abilities. We publicly release all metadata (including song names, links, and author attributions) and scraping code for academic research, allowing future studies to recreate the dataset, train, and benchmark for automatic transcription or other tasks.
@@ 3.1.2 Data Analysis
The dataset includes songs from over 4,000 unique artists, lasting on average 4.6 minutes/song. The average singing speed is found to be 90.1 Words Per Minute (WPM), less than half the Vietnamese spoken rate of 190 WPM [undefi]. The dataset includes various genres, as shown in Fig. 1. However, metadata extraction from zingmp3 alone resulted in many songs lacking genre information. Since Vietnamese music research lacks standardized genre classification tools or an agreed-upon genre ontology, Contrastive Learning of Musical Representations (CLMR) [undefj] was used to infer the top 50 MagnaTagATune tags [undefk]. The result was of weak alignment, with the 95th percentile confidence score reaching only 0.22. This suggests that this collection of Vietnamese music does not fit well within existing Western music’s tag ontology, highlighting the need for further research in Vietnamese music genre classification.
@@
Vietnamese dialects (which include both accents and variations in word choices, idioms, sentence structure, etc.) differ significantly across regions, a difference even more pronounced in singing, making accent classification valuable for music analysis. A pre-trained Wav2Vec2-base model [undefl, undefm] fine-tuned on Vietnamese Speech corpora [undefn, undefo] was used to predict gender [undefp] and dialect [undefq].
We extracted and ran inference on three 10-second vocal-only segments (isolated using Demucs v3 [undefr]) and determined the final classification through majority voting. For gender classification specifically, a “both” category was introduced for cases where at least one segment had a confidence score between 0.4 and 0.6, indicating ambiguity due to a unisex voice, absence of vocals, or a mix of male and female vocals. As shown in Fig. 2, the results indicate a distribution skewed towards Northern dialects, while Central dialects are underrepresented. Male-sung songs were more prevalent than those classified as “female” or “both”. This data imbalance is hard to address, as gender and accent metadata were unavailable during data collection. (Note that this observed skew can also stem from Wav2Vec2 model bias.)
@@ 3.2 Modeling
Since the current SOTA LyricWhiz model does not effectively handle Vietnamese songs, our approach is to fine-tune an ASR model end-to-end to serve as an ALT system. The Whisper model [undef] was chosen as the starting point, as it is renowned for its SOTA performance in multilingual tasks and exceptional fine-tuning capabilities.
Three Whisper model variants of different sizes (small, medium and large-v2) were fine-tuned and made available as checkpoints on Hugging Face [undefs].
@@ 3.2.1 Data Pre-processing
Audio segmentation of songs is performed using the provided ground truth timestamps from the song lyrics to chunk both the audio and lyrics into 30-second segments. This is required as Whisper processes audio in 30-second chunks [undef].
@@ 4.1 Finetuned Models
LyricWhiz was evaluated as a baseline for lyrics transcription, given its state-of-the-art performance in prior studies [undefb], though it had not been tested on Vietnamese songs. PhoWhisper [undeft], an ASR model fine-tuned on Vietnamese audio, was also assessed with the expectation of better performance on Vietnamese datasets. On VietLyrics test set, LyricWhiz produced a Word Error Rate (WER) of 49.22%, while PhoWhisper-large achieved 38.3%, shown in Table 3. However, both models fell short of the accuracy needed for effective Automatic Lyrics Transcription.
@@
We also experimented with training and evaluating only on vocals, separated using Demucs v3 [undefr]. However, the improvement observed after source separation was minimal for both models, with 1.3% WER improvement for whisper-medium and 0.5% WER improvement for whisper-large-v2. Given the negligible benefits and the added complexity of these extra pre-processing steps, the decision was made to exclude these steps from the final approach.
Lower-case evaluation was also found to produce better performance metrics compared to case-sensitive evaluation, with WER decreasing by up to 5% on large, as shown in Table 4. This intuitive result suggests that relying on the LLM-based transcriber for capitalization may not be the optimal approach. Instead, more effective methods for inferring line breaks could involve post-processing techniques, such as detecting silence between singing segments or leveraging the approaches introduced in Cífka et al. [undefu].
@@
Model WER (Case-sensitive) WER (Lowercase)
@@
Since hallucination from non-vocal audio is a major challenge in ASR models, experiments were conducted to suppress <nospeech> tokens. Following the original Whisper paper [undef], tokens were filtered when the <nospeech> probability exceeded 0.6 and the log confidence probability was below -1. These experiments were evaluated on 100 randomly sampled songs from the 1,000-song validation set. PhoWhisper was tested under various configurations, with vocals-only tracks extracted using Demucs. As shown in Table 5, for PhoWhisper-small, combining source-separated audio with <nospeech> suppression provided the best performance. However, for Whisper large, testing on raw audio without suppression performed best, with our fine-tuned Whisper significantly outperforming others.
This outcome is consistent with the findings in Cífka et al. [undefu], where the transcriptions from only vocals are less accurate than those from raw data for large models. This is preferable, as it enables deploying an end-to-end model without dependence on source separation performance. Furthermore, qualitative reviews of our finetuned Whisper’s inferences found no instances of long hallucinations, resulting in the discontinuation of further silence suppression experiments.
@@ 5 Conclusion
We collected VietLyrics, a large-scale diverse dataset of Vietnamese music with lyrics, and fine-tuned the Whisper models [undef] to develop an Automatic Lyrics Transcription system. Our best performing model, Whisper-large-v2, achieved a WER of 24.61% for case-sensitive lyrics and 20.52% for lowercase lyrics, demonstrating its effectiveness in transcribing Vietnamese songs of various dialects, genres and genders. We publicly release the VietLyrics dataset [undefv] and models [undefw, undefx, undefy], hoping that this work inspires further research and improvements in Automatic Lyrics Transcription and music computing in general, for the Vietnamese language and music.